# Learning Interactive Real-World Simulators

### v2

@@ Abstract
Generative models trained on internet data have revolutionized how text, image, and video content can be created. Perhaps the next milestone for generative models is to simulate realistic experience in response to actions taken by humans, robots, and other interactive agents. Applications of a real-world simulator range from controllable content creation in games and movies, to training embodied agents purely in simulation that can be directly deployed in the real world. We explore the possibility of learning a universal simulator of real-world interaction through generative modeling. We first make the important observation that natural datasets available for learning a real-world simulator are often rich along different dimensions (e.g., abundant objects in image data, densely sampled actions in robotics data, and diverse movements in navigation data). With careful orchestration of diverse datasets, each providing a different aspect of the overall experience, we can simulate the visual outcome of both high-level instructions such as “open the drawer” and low-level controls such as “move by $\Delta x,\Delta y$” from otherwise static scenes and objects. We use the simulator to train both high-level vision-language policies and low-level reinforcement learning policies, each of which can be deployed in the real world in zero shot after training purely in simulation. We also show that other types of intelligence such as video captioning models can benefit from training with simulated experience, opening up even wider applications. Video demos can be found at universal-simulator.github.io.
@@ 1 Introduction
Generative models trained on internet data can now produce highly realistic text (openai2023gpt4, ),image (ramesh2022hierarchical, ), and video (ho2022imagen, ). Perhaps the ultimate goal of generative models is to be able to simulate the visual effects of a wide variety of actions, from how cars are driven on a street to how furniture and meals are prepared. With a real-world simulator, humans can “interact” with diverse scenes and objects, robots can learn from simulated experience without risking physical damage, and a vast amount of “real-world” like data can be simulated to train other types of machine intelligence.
@@
One roadblock to building this simulator lies in the datasets — different datasets cover different information that have to be brought together to simulate realistic experience. For instance, paired text-image data from the internet contains rich scenes and objects but little movement (schuhmann2022laion, ; zhai2022scaling, ), video captioning and question answering data contain rich high-level descriptions but little low-level movement detail (xu2016msr, ; krishna2017dense, ), human activity data contains rich human action but little mechanical motion (miech2019howto100m, ; grauman2022ego4d, ), and robotics data contains rich robot action but are limited in quantity (dasari2019robonet, ; mandlekar2018roboturk, ). Since different datasets are curated by different industrial or research communities for different purposes, divergence in information is natural and hard to overcome, posing difficulties to a real-world simulator that seeks to capture all visual aspects of the world.
@@
In this work, we propose to combine a wealth of data in a conditional video generation framework to instantiate a universal simulator [^1] Note that by “universal”, we mean the model can simulate through the unified interface of actions and videos, as opposed to being able to simulate everything. Sound, for instance, is not being simulated.. Under a unified action-in-video-out interface, the simulator enables rich interaction through fine-grained motion control of otherwise static scenes and objects. To support long-horizon repeated interactions, we formulate the simulator as an observation prediction model that can be rolled out autoregressively to support consistent simulation across video generation boundaries.
@@
While the potential applications of the simulator are broad, we demonstrate three specific use cases. We first show how the simulator enables a vision-language policy to perform long-horizon goal-conditioned tasks through hindsight relabeling of simulated experience (andrychowicz2017hindsight, ). In addition to learning high-level vision-language policies, we illustrate how the simulator can enable learning low-level control policies by leveraging model-based reinforcement learning (RL) (sutton1988learning, ). Both the high-level vision-language policy and the low-level control policy, while trained purely in simulation, can generalize to real robot settings. This is enabled by using the simulator that is nearly visually indistinguishable from the real world, achieving one step towards bridging the sim-to-real gap in embodied learning (rusu2017sim, ). Furthermore, we can simulate rare events where data collection is expensive or dangerous (e.g., crashes in self-driving cars). Such simulated videos can then be used to improve other machine intelligence such as rare event detectors, suggesting broad applications of a universal simulator beyond embodied learning. The main contributions can be summarized as follows:
@@
Figure 1: A universal simulator. The simulator of the real-world learns from broad data with diverse information including objects, scenes, human activities, motions in navigation and manipulation, panorama scans, and simulations and renderings.
@@ 2 Learning an Interactive Real-World Simulator
Figure 2: Training and inference of the UniSim. The UniSim is a video diffusion model trained to predict the next (variable length) set of observation frames ($o_{t}$) given observations from the past (e.g., $o_{t-1}$) and action input $a_{t-1}$. The UniSim can handle temporally extended actions in various modalities such as motor controls ($\Delta x_{1},\Delta\omega_{1},\Delta x_{2},...$), language descriptions (“wipe table”), and actions extracted from camera motions and other sources. Each dotted arrow indicates concatenating the initial noise sample for the next video segment with the previous frame.
@@ 2.1 Orchestrating Diverse Datasets
Below, we highlight diverse information in different datasets and propose ways to process actions into a common format (see all datasets used to train the UniSim in Appendix 9).
  • Simulated execution and renderings. While annotating actions for real-world videos is expensive, simulation engines such as Habitat (savva2019habitat, ) can render a wide variety of actions. We use datasets previously collected from these simulators, i.e., Habitat object navigation with HM3D (ramakrishnan2021hm3d, ) and Language Table Data from lynch2023interactive to train the UniSim. We extract text descriptions as actions when available. For simulated continuous control actions, we encode them via language embeddings and concatenate the text embeddings with discretized control values.
  • Real robot data. An increasing amount of video data of real-robot executions paired with task descriptions such as the Bridge Data (ebert2021bridge, ) and data that enabled RT-1 and RT-2 (brohan2022rt, ) are becoming increasingly available. Despite low-level control actions often being different across robots, the task descriptions can serve as high-level actions in the UniSim. We further include discretize continuous controls actions when available similar to simulated robotics data.
  • Human activity videos. Rich human activity data such as Ego4D (grauman2022ego4d, ), EPIC-KITCHENS (damen2018scaling, ), and Something-Something V2 (goyal2017something, ) have been curated. Different from low-level robot controls, these activities are high-level actions that humans take to interact with the world. But these actions are often provided as labels for video classification or activity recognition tasks (goyal2017something, ). In this case, we convert the video labels into text actions. In addition, we subsample the videos to construct chunks of observations at a frame rate that captures meaningful actions.
  • Panorama scans. There exists a wealth of 3D scans such as Matterport3D (chang2017matterport3d, ). These static scans do not contain actions. We construct actions (e.g., turn left) by truncating panorama scans and utilize the information of camera poses between two images.
  • Internet text-image data. Paired text-image datasets such as LAION (schuhmann2021laion, ) contain static images of a variety of objects without actions. However, the captions often contain motion information such as “a person walking”. To use image data in the UniSim, we treat individual images as single-frame videos and image captions as actions.
For each of these datasets, we process text tokens into continuous representations using T5 language model embeddings (raffel2020exploring, ) concatenated with low-level actions such as robot controls. This serves as the final unified action space of our simulator.
@@ 2.2 Simulating Long-Horizon Interactions through Observation Prediction
Simulating Real-World Interactions. We define an observation space $\mathcal{O}$ and an action space $A$ which capture the videos and actions described in Section 2.1. At a specific interactive step $t$, an agent, having observed a set of history frames $h_{t-1}\in O$, decides on some temporally extended action $a_{t-1}\in A$, which can be resolved into a sequence of low-level robot commands to be executed in the real world. During the execution, the next set of video frames $o_{t}\in O$ are captured from the real world. The goal of a simulator is to predict $o_{t}$ from $h_{t-1}$ and $a_{t-1}$. We can formulate this prediction problem as learning an observation prediction model $p(o_{t}|h_{t-1},a_{t-1})$. While an ideal predictive model should condition on all information of the past, i.e., $(o_{0},a_{0}\ldots,a_{t-2},o_{t-1})$, through some recurrent state, we found conditioning on a finite set of frames (e.g., frames from the most recent interaction, $o_{t-1}$) greatly simplifies the modeling problem. To simulate long interactions, we can sample from the observation prediction model $p(o_{t}|h_{t-1},a_{t-1})$ autoregressively conditioned on the previously sampled observations. One advantage of this observation prediction model is that the simulator stays the same across all tasks and can be used in combination with any reward function, which can be separately learned. The learned reward function can then be used to optimize policies $\pi(a_{t}|h_{t})$ using existing decision making algorithms such as planning and RL, as we will illustrate in Section 4.1 and Section 4.2.
Parametrizing and Training the Simulator. We parametrize $p(o_{t}|h_{t-1},a_{t-1})$ using diffusion models (sohl2015deep, ; ho2020denoising, ) as an instantiation of the UniSim outlined in Figure 2. Specifically, the reverse process learns a denoising model $\epsilon_{\theta}(o_{t}^{(k)},k|h_{t-1},a_{t-1})$ that, conditioned on the history, generates the next observationfrom initial noise samples using $K$ denoising steps. In practice, we only use previous video frames and omit previous actions as history, and concatenate previous video frames with initial noise samples $o^{(K)}_{t}\sim\mathcal{N}(0,I)$ channelwise to serve as conditional inputs to the denoising model. To condition on an action $a_{t-1}$, we leverage classifier-free guidance (ho2022classifier, ). The final $\overline{\mathcal{T}}(o_{t}|h_{t-1},a_{t-1})$ is parametrized by the variance schedule:
@@
Architecture and Training. We use the video U-Net architecture (ho2022video, ) to implement the UniSim by employing interleaved temporal and spatial attention and convolution layers in both the downsampling and upsampling passes. For history conditioning, we replicate the conditioning frames at all future frame indices, and concatenate the conditioning frames with the noise sample for each of the future frame to serve as input to the U-Net. The UniSim model has 5.6B parameters and requires 512 TPU-v3 and 20 days to train on all data. See more details in Appendix 10.
@@ 3 Simulating Real-World Interactions
Figure 3: Action-rich simulations. The UniSim can support manipulation actions such as “cut carrots”, “wash hands”, and “pickup bowl” from the same initial frame (top left) and other navigation actions.
@@ 3.1 Action-Rich, Long-Horizon, and Diverse Interactions
Action-Rich Simulation. We first demonstrate action-rich interactions through natural language actions. Figure 3 shows simulation of human manipulation and navigation starting from the same initial observation (left-most column). We can instruct a person in the initial frame to perform various kitchen tasks (top left), press different switches (top right), or navigate scenes (bottom). The model only trained on generic internet data, without action-rich manipulation data such as EPIC-KITCHENS (damen2018scaling, ), fails to simulate action-rich manipulations (Appendix 13).
Figure 4: Long-horizon simulations. The UniSim sequentially simulates 8 interactions autoregressively. The simulated interactions maintain temporal consistency across long-horizon interactions, correctly preserving objects and locations (can on counter in column 2-7, orange in drawer in column 4-5).
Long-Horizon Simulation. Next, we illustrate 8 sequential interactions in Figure 4. We condition the simulation of each interaction on previous observations and new language action as described in Section 2.2. The UniSim successfully preserves objects manipulated by previous instructions (e.g., the orange and can are preserved in the drawers in Columns 4, 5, 7, 8 after being put in the drawers). See additional long-horizon interactions in Appendix 8.1.
Diversity and Stochasticity in the Simulator. The UniSim can also support highly diverse and stochastic environment transitions, e.g., diverse objects being revealed after removing the towel on top (Figure 5 left), diverse object colors and locations (cups and pens in Figure 5 right), and real-world variabilities such as change in camera angles. Flexibility in diffusion models promotes simulation of highly stochastic environments that cannot be controlled by actions, so that a policy can learn to only control the controllable part (yang2022dichotomy, ).
@@ 3.2 Ablation and Analysis
Condition FID $\downarrow$ FVD $\downarrow$ IS $\uparrow$ CLIP $\uparrow$ 1 frame 59.47 315.69 3.03 22.55 4 distant 34.89 237 3.43 22.62 4 recent 34.63 211.3 3.52 22.63
[Uncaptioned image] Figure 6: Simulations of low-data domains using the Habitat object navigation using HM3D dataset (ramakrishnan2021hm3d, ) with only 700 training examples. Prefixing language actions with dataset identifier leads to video samples that complete the action (top).
Frame Conditioning Ablations. We ablate over choices of past frames to condition on using a validation split of the Ego4D dataset (grauman2022ego4d, ), which contains egocentric movement requiring proper handling of observation history. We compare the UniSim conditioned on different numbers of past frames in Table 1. Conditioning on 4 frames is better than conditioning on a single frame, but conditioning on history that is too far in the past (4 frames with exponentially increasing distances) can hurt performance. Increasing the number of conditioning frames beyond 4 did not further improve performance on Ego4D, but it could be helpful for applications that require memory from distant past (e.g., navigation for retrieval).
Simulating Low-Data Domains. During joint training of the UniSim on diverse data, we found that naïvely combining datasets of highly varying size can result in low generation quality in low-data domains. While we can increase the weight of these domains in the data mixture during training, we found that attaching a domain identifier such as the name of the dataset to the actions being conditioned on improves generation quality in low-data domains, as shown in Figure 6. While such domain identifier improves in-distribution generation quality, we found domain-specific identifiers to hurt generalization to other domains, and should only be applied with the test domain is in distribution of the training domain.
4 Applications of the UniSim
We now demonstrate how the UniSim can be used to train other types of machine intelligence such as vision-language policies, RL agents, and vision-language models through simulating highly realistic experiences.

### v3

@@ Abstract
Generative models trained on internet data have revolutionized how text, image, and video content can be created. Perhaps the next milestone for generative models is to simulate realistic experience in response to actions taken by humans, robots, and other interactive agents. Applications of a real-world simulator range from controllable content creation in games and movies, to training embodied agents purely in simulation that can be directly deployed in the real world. We explore the possibility of learning a universal simulator (UniSim) of real-world interaction through generative modeling. We first make the important observation that natural datasets available for learning a real-world simulator are often rich along different dimensions (e.g., abundant objects in image data, densely sampled actions in robotics data, and diverse movements in navigation data). With careful orchestration of diverse datasets, each providing a different aspect of the overall experience, we can simulate the visual outcome of both high-level instructions such as “open the drawer” and low-level controls such as “move by $\Delta x,\Delta y$” from otherwise static scenes and objects. We use the simulator to train both high-level vision-language policies and low-level reinforcement learning policies, each of which can be deployed in the real world in zero shot after training purely in simulation. We also show that other types of intelligence such as video captioning models can benefit from training with simulated experience, opening up even wider applications. Video demos can be found at https://universal-simulator.github.io.
@@ 1 Introduction
Generative models trained on internet data can now produce highly realistic text (OpenAI, 2023),image (Ramesh et al., 2022), and video (Ho et al., 2022a). Perhaps the ultimate goal of generative models is to be able to simulate the visual effects of a wide variety of actions, from how cars are driven on a street to how furniture and meals are prepared. With a real-world simulator, humans can “interact” with diverse scenes and objects, robots can learn from simulated experience without risking physical damage, and a vast amount of “real-world” like data can be simulated to train other types of machine intelligence.
@@
One roadblock to building this simulator lies in the datasets — different datasets cover different information that have to be brought together to simulate realistic experience. For instance, paired text-image data from the internet contains rich scenes and objects but little movement (Schuhmann et al., 2022; Zhai et al., 2022), video captioning and question answering data contain rich high-level descriptions but little low-level movement detail (Xu et al., 2016; Krishna et al., 2017), human activity data contains rich human action but little mechanical motion (Miech et al., 2019; Grauman et al., 2022), and robotics data contains rich robot action but are limited in quantity (Dasari et al., 2019; Mandlekar et al., 2018). Since different datasets are curated by different industrial or research communities for different purposes, divergence in information is natural and hard to overcome, posing difficulties to a real-world simulator that seeks to capture all visual aspects of the world.
@@
In this work, we propose to combine a wealth of data in a conditional video generation framework to instantiate a universal simulator (UniSim) [^1] Note that by “universal”, we mean the model can simulate through the unified interface of actions and videos, as opposed to being able to simulate everything. Sound, for instance, is not being simulated.. Under a unified action-in-video-out interface, the simulator enables rich interaction through fine-grained motion control of otherwise static scenes and objects. To support long-horizon repeated interactions, we formulate the simulator as an observation prediction model that can be rolled out autoregressively to support consistent simulation across video generation boundaries.
@@
While the potential applications of the simulator are broad, we demonstrate three specific use cases. We first show how the simulator enables a vision-language policy to perform long-horizon goal-conditioned tasks through hindsight relabeling of simulated experience (Andrychowicz et al., 2017). In addition to learning high-level vision-language policies, we illustrate how the simulator can enable learning low-level control policies by leveraging model-based reinforcement learning (RL) (Sutton, 1988). Both the high-level vision-language policy and the low-level control policy, while trained purely in simulation, can generalize to real robot settings. This is enabled by using the simulator that is nearly visually indistinguishable from the real world, achieving one step towards bridging the sim-to-real gap in embodied learning (Rusu et al., 2017). Furthermore, we can simulate rare events where data collection is expensive or dangerous (e.g., crashes in self-driving cars). Such simulated videos can then be used to improve other machine intelligence such as rare event detectors, suggesting broad applications of UniSim beyond embodied learning. The main contributions can be summarized as follows:
@@
Figure 1: A universal simulator (UniSim). The simulator of the real-world learns from broad data with diverse information including objects, scenes, human activities, motions in navigation and manipulation, panorama scans, and simulations and renderings.
@@ 2 Learning an Interactive Real-World Simulator
Figure 2: Training and inference of UniSim. UniSim is a video diffusion model trained to predict the next (variable length) set of observation frames ($o_{t}$) given observations from the past (e.g., $o_{t-1}$) and action input $a_{t-1}$. UniSim can handle temporally extended actions in various modalities such as motor controls ($\Delta x_{1},\Delta\omega_{1},\Delta x_{2},...$), language descriptions (“wipe table”), and actions extracted from camera motions and other sources. Each dotted arrow indicates concatenating the initial noise sample for the next video segment with the previous frame.
@@ 2.1 Orchestrating Diverse Datasets
Below, we highlight diverse information in different datasets and propose ways to process actions into a common format (see all datasets used to train UniSim in Appendix B).
  • Simulated execution and renderings. While annotating actions for real-world videos is expensive, simulation engines such as Habitat (Savva et al., 2019) can render a wide variety of actions. We use datasets previously collected from these simulators, i.e., Habitat object navigation with HM3D (Ramakrishnan et al., 2021) and Language Table Data from Lynch et al. (2023) to train UniSim. We extract text descriptions as actions when available. For simulated continuous control actions, we encode them via language embeddings and concatenate the text embeddings with discretized control values.
  • Real robot data. An increasing amount of video data of real-robot executions paired with task descriptions such as the Bridge Data (Ebert et al., 2021) and data that enabled RT-1 and RT-2 (Brohan et al., 2022) are becoming increasingly available. Despite low-level control actions often being different across robots, the task descriptions can serve as high-level actions in UniSim. We further include discretize continuous controls actions when available similar to simulated robotics data.
  • Human activity videos. Rich human activity data such as Ego4D (Grauman et al., 2022), EPIC-KITCHENS (Damen et al., 2018), and Something-Something V2 (Goyal et al., 2017) have been curated. Different from low-level robot controls, these activities are high-level actions that humans take to interact with the world. But these actions are often provided as labels for video classification or activity recognition tasks (Goyal et al., 2017). In this case, we convert the video labels into text actions. In addition, we subsample the videos to construct chunks of observations at a frame rate that captures meaningful actions.
  • Panorama scans. There exists a wealth of 3D scans such as Matterport3D (Chang et al., 2017). These static scans do not contain actions. We construct actions (e.g., turn left) by truncating panorama scans and utilize the information of camera poses between two images.
  • Internet text-image data. Paired text-image datasets such as LAION (Schuhmann et al., 2021) contain static images of a variety of objects without actions. However, the captions often contain motion information such as “a person walking”. To use image data in UniSim, we treat individual images as single-frame videos and image captions as actions.
For each of these datasets, we process text tokens into continuous representations using T5 language model embeddings (Raffel et al., 2020) concatenated with low-level actions such as robot controls. This serves as the final unified action space of our simulator.
@@ 2.2 Simulating Long-Horizon Interactions through Observation Prediction
Simulating Real-World Interactions. We define an observation space $O$ and an action space $A$ which capture the videos and actions described in Section 2.1. At a specific interactive step $t$, an agent, having observed a set of history frames $h_{t-1}\in O$, decides on some temporally extended action $a_{t-1}\in A$, which can be resolved into a sequence of low-level robot commands to be executed in the real world. During the execution, the next set of video frames $o_{t}\in O$ are captured from the real world. The goal of a simulator is to predict $o_{t}$ from $h_{t-1}$ and $a_{t-1}$. We can formulate this prediction problem as learning an observation prediction model $p(o_{t}|h_{t-1},a_{t-1})$. While an ideal predictive model should condition on all information of the past, i.e., $(o_{0},a_{0}\ldots,a_{t-2},o_{t-1})$, through some recurrent state, we found conditioning on a finite set of frames (e.g., frames from the most recent interaction, $o_{t-1}$) greatly simplifies the modeling problem. To simulate long interactions, we can sample from the observation prediction model $p(o_{t}|h_{t-1},a_{t-1})$ autoregressively conditioned on the previously sampled observations. One advantage of this observation prediction model is that the simulator stays the same across all tasks and can be used in combination with any reward function, which can be separately learned. The learned reward function can then be used to optimize policies $\pi(a_{t}|h_{t})$ using existing decision making algorithms such as planning and RL, as we will illustrate in Section 4.1 and Section 4.2.
Parametrizing and Training the Simulator. We parametrize $p(o_{t}|h_{t-1},a_{t-1})$ using diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) as an instantiation of UniSim outlined in Figure 2. Specifically, the reverse process learns a denoising model $\epsilon_{\theta}(o_{t}^{(k)},k|h_{t-1},a_{t-1})$ that, conditioned on the history, generates the next observationfrom initial noise samples using $K$ denoising steps. In practice, we only use previous video frames and omit previous actions as history, and concatenate previous video frames with initial noise samples $o^{(K)}_{t}\sim\mathcal{N}(0,I)$ channelwise to serve as conditional inputs to the denoising model. To condition on an action $a_{t-1}$, we leverage classifier-free guidance (Ho & Salimans, 2022). The final $\overline{\mathcal{T}}(o_{t}|h_{t-1},a_{t-1})$ is parametrized by the variance schedule:
@@
Architecture and Training. We use the video U-Net architecture (Ho et al., 2022b) to implement UniSim by employing interleaved temporal and spatial attention and convolution layers in both the downsampling and upsampling passes. For history conditioning, we replicate the conditioning frames at all future frame indices, and concatenate the conditioning frames with the noise sample for each of the future frame to serve as input to the U-Net. UniSim model has 5.6B parameters and requires 512 TPU-v3 and 20 days to train on all data. See more details in Appendix C.
@@ 3 Simulating Real-World Interactions
Figure 3: Action-rich simulations. UniSim can support manipulation actions such as “cut carrots”, “wash hands”, and “pickup bowl” from the same initial frame (top left) and other navigation actions.
@@ 3.1 Action-Rich, Long-Horizon, and Diverse Interactions
Action-Rich Simulation. We first demonstrate action-rich interactions through natural language actions. Figure 3 shows simulation of human manipulation and navigation starting from the same initial observation (left-most column). We can instruct a person in the initial frame to perform various kitchen tasks (top left), press different switches (top right), or navigate scenes (bottom). The model only trained on generic internet data, without action-rich manipulation data such as EPIC-KITCHENS (Damen et al., 2018), fails to simulate action-rich manipulations (Appendix F).
Figure 4: Long-horizon simulations. UniSim sequentially simulates 8 interactions autoregressively. The simulated interactions maintain temporal consistency across long-horizon interactions, correctly preserving objects and locations (can on counter in column 2-7, orange in drawer in column 4-5).
Long-Horizon Simulation. Next, we illustrate 8 sequential interactions in Figure 4. We condition the simulation of each interaction on previous observations and new language action as described in Section 2.2. UniSim successfully preserves objects manipulated by previous instructions (e.g., the orange and can are preserved in the drawers in Columns 4, 5, 7, 8 after being put in the drawers). See additional long-horizon interactions in Appendix A.1.
Diversity and Stochasticity in the Simulator. UniSim can also support highly diverse and stochastic environment transitions, e.g., diverse objects being revealed after removing the towel on top (Figure 5 left), diverse object colors and locations (cups and pens in Figure 5 right), and real-world variabilities such as change in camera angles. Flexibility in diffusion models promotes simulation of highly stochastic environments that cannot be controlled by actions, so that a policy can learn to only control the controllable part (Yang et al., 2022).
@@ 3.2 Ablation and Analysis
Condition FID $\downarrow$ FVD $\downarrow$ IS $\uparrow$ CLIP $\uparrow$
1 frame 59.47 315.69 3.03 22.55
4 distant 34.89 237 3.43 22.62
4 recent 34.63 211.3 3.52 22.63
@@
Figure 6: Simulations of low-data domains using the Habitat object navigation using HM3D dataset (Ramakrishnan et al., 2021) with only 700 training examples. Prefixing language actions with dataset identifier leads to video samples that complete the action (top).
Frame Conditioning Ablations. We ablate over choices of past frames to condition on using a validation split of the Ego4D dataset (Grauman et al., 2022), which contains egocentric movement requiring proper handling of observation history. We compare UniSim conditioned on different numbers of past frames in Table 1. Conditioning on 4 frames is better than conditioning on a single frame, but conditioning on history that is too far in the past (4 frames with exponentially increasing distances) can hurt performance. Increasing the number of conditioning frames beyond 4 did not further improve performance on Ego4D, but it could be helpful for applications that require memory from distant past (e.g., navigation for retrieval).
Simulating Low-Data Domains. During joint training of UniSim on diverse data, we found that naïvely combining datasets of highly varying size can result in low generation quality in low-data domains. While we can increase the weight of these domains in the data mixture during training, we found that attaching a domain identifier such as the name of the dataset to the actions being conditioned on improves generation quality in low-data domains, as shown in Figure 6. While such domain identifier improves in-distribution generation quality, we found domain-specific identifiers to hurt generalization to other domains, and should only be applied with the test domain is in distribution of the training domain.
4 Applications of UniSim
We now demonstrate how UniSim can be used to train other types of machine intelligence such as vision-language policies, RL agents, and vision-language models through simulating highly realistic experiences.